OpenStudy (samara8954):

A plane can fly 66 miles in 5 minutes. At this rate, how many miles can the plane fly in 18 minutes? 244.2 224.4 231 237.6

OpenStudy (samara8954):

@aaliyahfairgood

jimthompson5910 (jim_thompson5910):

(# of miles)/(# of minutes) = (# of miles)/(# of minutes) (66 miles)/(5 minutes) = (x miles)/(18 minutes) 66/5 = x/18 66*18 = 5*x ... cross multiply x = ???

OpenStudy (samara8954):

can you help me with my next problem @aaliyahfairgood

OpenStudy (aaliyahfairgood):

Yes